OpenStudy (anonymous):

y=-3/2x+4 Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

the graph intercepts the y-axis at y=4. so the point (0,4) would be marked. form (0,4), go to the left 2 units and up 3units to (-2, 7) draw a line through these 2 points

OpenStudy (anonymous):

Can you help me with another one? please.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

–3x + 2y = 8 Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

slope intercept form is y=mx+b so your equation goes to 2y= 3x+8 y=3/2x+4

OpenStudy (anonymous):

Slope= 3/2 Y-intercept= 4

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Wait, how would you graph it?, lol

OpenStudy (anonymous):

I thought you where asking me for the slope and Y-intercept.

OpenStudy (anonymous):

y intercept is 4. So (0,4) is plotted. then go to the right 2 units and up 3 units to (2,7) draw a line through the 2 points

OpenStudy (anonymous):

In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences